Find the margin of error for the given values of​ c, sigma​, and n. c=0.95​, sigma=3.9​, n=100
Answer #1

Solution :

Given that,

Sample size = n = 100

Z\alpha/2 = 1.96

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 1.96 * (3.9 / \sqrt 100)

Margin of error = E = 0.8